(2.1) Find parametric equations of the line that passes through the point P = (2, 0, −1) and is

parallel to the vector ~n =< 2, 1, 3 >.

(2.2) Find parametric equations of the line that passes through the points A = (1, 2, −3) and B =

(7, 2, −4).

(2.3) Find parametric equations for the line of intersection of the planes −5x + y − 2z = 3 and

2x − 3y + 5z = −7.


1
Expert's answer
2021-07-01T15:46:14-0400

Vector equation of a line is:

"r=r_0+tv"


where r0 is position vector of any point on the line.

And v is a vector parallel to the line.


2.1

it is given that (2,0,-1) is a point on the line.

Its position vector is "r_0=2i+0j-k"

Therefore the vector equation of the line is

"r=(2i+oj-k)+t(2i+j+3k)\\r=(2+2t)i+tj+(-1+3t)k"


Therefore the parametric equations are

"x=2+2t, \space y=t, \space z=-1+3t"



2.2

it is given that (1,2,-3) and (7,2,-4) are points on the line.


The direction vector (a,b,c) will be"<7-1,2-2,-4-(-3)>\\<a,b,c>=<6,0,-1>"

The vector equation of the line is


"r=(x_0,y_0,z_0)+t(a,b,c)\\r=(1,2,-3)+t(6,0,-1)"


Therefore the parametric equations are

"x=1+6t, \space y=2, \space z=-3-t"


2.3

"-5x+y-2z=3\\2x-3y+5z=-7"


set z=0, solve for x and y

"-5x+y-0=3\\2x-3y+0=-7"


"y=5x+3\\2x-3(5x+3)=-7\\2x-15x-9=-7\\-13x=2\\x=-\frac{2}{13}"


"y=5(-\frac{2}{13})+3\\=-\frac{10}{13}+3\\=\frac{29}{13}"


point 1: "(x_0,y_0,z_0)=(-\frac{2}{13},\frac{29}{13},0)"

We can use the cross product to find a direction vector for the line


The cross product of the normals to the planes results in a vector that is orthogonal to both and would be parallel to the line of intersection.


"<a,b,c>\times<d,e,f>=<bf-ce, cd-af, ae-bd>\\<-5,1,-2>\times<2,-3,5>=1(5)--2(-3),-2(2)--5(5),-5(-3)-1(2)\\=<-1,21,13>"


plug the point into "(x_0,y_0,z_0)" and the vector found into <a,b,c>


"x=-\frac{2}{13}-t"


"y=\frac{29}{13}+21t"


"z=13t"